1. The little Thrifty Market has a parking lot with dimensions 125 feet by 210 feet. How many square yard is that?
2. The posted speed limit on many U.S highways is 65 miles per hour. What is this speed in feet per seconds?

Answer by mananth(16949)

The little Thrifty Market has a parking lot with dimensions 125 feet by 210 feet. How many square yard is that?

3feet make an yard
125 feet = 125/3 = 41.66 yards
210 feet = 210/3 = 70 yards
Area = L * W
41.66 * 70 = 2916.66 square yards
2. The posted speed limit on many U.S highways is 65 miles per hour. What is this speed in feet per seconds?
to convert mph to feet per second
1 mile = 5280 feet
1 hour = 3600 seconds
mile/hour = 5280 feet /3600 seconds
1 mile / hour = 1.466 feet /second
65 mph = 1.466* 65 = 95.33 feet/second